# Simulations in Statistical Workflows

### v1

@@ 1 Introduction
In statistics, the emergence of simulation is typically associated with the birth of Monte Carlo methods in Los Alamos during World War II (Robert and Casella,, 2011), eventually leading to the publication of the Metropolis algorithm in the early 1950s (Metropolis et al.,, 1953). Hastings, (1970) and Peskun, (1973) later generalized the Metropolis algorithm as a family of simulation-based tools, namely, Markov chain Monte Carlo (MCMC), designed to mitigate the curse of dimensionality in direct Monte Carlo estimation. Crucially, the inception of Monte Carlo methods coincides with the end of what Efron and Hastie, (2021) characterize as classical statistical inference and the beginning of early computer-age methods, such as bootstrap and jackknife, that demand randomized resampling and mechanical repetition.
@@
Early MCMC methods were mainly confined to analytically tractable statistical models, defined through closed-form distributions (e.g., Gaussian). In contrast, implicit models—defined through a stochastic mechanism (e.g., a randomized simulation) without a known distribution for computing the likelihood of their outputs (Diggle and Gratton,, 1984)—remained out of reach for Bayesian methods until the advent of approximate Bayesian computation (ABC; Rubin,, 1984; Tavaré et al.,, 1997), which leverages simulations to approximate Bayesian posteriors. In ABC, simulation explicitly takes on a dual role: it serves both as an epistemic tool for solving forward problems (from unknowns to knowns) and as a computational tool for solving inverse problems (from knowns to unknowns). Indeed, the coinage of the term simulation-based inference (SBI; Cranmer et al.,, 2020) is reflective of this role.
Figure 1: Forward simulation and inverse inference are two central components in a probabilistic joint model of unknowns $\theta$ and observables $y$.
As the notion of statistical model loses its well-defined contours in an AI-dominated conceptual landscape, the role of simulation in statistics becomes increasingly difficult to delineate from its role in science in general. However, since this paper focuses on statistical workflows—iterative processes of model building and model criticism—our discussion will center on the versatility of simulating from statistical models. In particular, this means generating random draws from model-implied distributions, much like drawing random cards from a cleverly arranged deck. The organization of the following sections elucidates use cases for such random draws throughout the key stages of a minimalist statistical workflow, namely: 1) model specification; 2) model verification; 3) model inference; and 4) model checking. We conclude with an outlook on simulation intelligence and the ascent of automated statistics.
@@ 2 Model Specification
  • P (Probability Distribution) – the probabilistic joint model that defines the relationships between model parameters, latent variables, and observed data.
@@
According to the PAD model taxonomy, model specification refers to defining the components of a P model. In a Bayesian settings, this entails formulating a prior $\pi(\theta)$ and an observation model $\pi(y\mid\theta)$ which together determine the generative joint model $\pi(\theta,y)$. In a frequentist settings, the spotlight falls on the observation model $\pi(y\mid\theta)$ which is defined over a specific domain $\Theta\ni\theta$, and the P model reduces to the observation model alone.
As part of a Bayesian workflow, the plausibility of P models can already be evaluated through prior predictive or prior pushforward checks (Gelman et al.,, 2020). The purpose of these checks is to assess whether the generative behavior of a P model is consistent with the available domain expertise via simulations from the model-implied (i.e., prior predictive) distribution, $\pi(y)=\int\pi(y\mid\theta)\pi(\theta)\mathrm{d}\theta$. Consistency with domain expertise not only figures in domain-specific Bayesian workflows (Schad et al.,, 2021), but can also be considered as a crucial aspect of theory building in general.
Here, once again, simulation serves as an antidote to complexity: consistency with domain expertise may not be immediately evident from the model formulation alone, even when a model’s specification appears beguilingly simple (see, for example, Aguilera-Venegas et al.,, 2019, for a probabilistic extension of Conway’s Game of Life). Moreover, if the data $y$ is high-dimensional, the success of consistency checks can be challenging to assess. This is where prior pushforward checks come into play, following the procedure suggested by Schad et al., (2021): (1) define a low-dimensional, interpretable summary statistic $T(y)$; (2) determine plausible value regions for the summary statistic; and (3) simulate model predictions to ensure the summary statistic falls within these regions [^1] In a frequentist setting, the prior can be replaced with a set of plausible values for the parameters $\Theta_{\text{plausible}}$..
@@
Figure 2: Graphical illustration of the simulation-based framework for prior elicitation proposed by Bockting et al., 2024b . The process starts by identifying target quantities from the domain expert and selecting elicitation techniques to obtain expert-elicited statistics. Predictions are then simulated by sampling from the prior $p_{\lambda}(\theta)$ and computing model-implied target quantities. A loss function $L_{m}$ assesses the consistency between model and expert-elicited statistics and adjusts hyperparameters $\lambda$ to minimize discrepancies. The process continues until the prior aligns closely with expert knowledge.
@@ 3 Model Verification
Before deploying a statistical model on real data, we must first ensure that the interplay between the probabilistic joint model P and the approximator A functions as intended in the scenarios they were designed for, namely, when their underlying assumptions hold. In many cases, simulations offer the only feasible means of verifying such correctness, or at least gauging in silico whether inferences are sufficiently informative for their intended purposes.
@@ 3.1 Verifying Calibration
Correct uncertainty calibration is arguably one of the most important properties of any method estimating quantities from statistical models fitted to data (Little,, 2006; Gneiting et al.,, 2007). Informally, we can say that the uncertainty of an approximator is well calibrated if it captures the true amount of uncertainty in a system. Except for simple special cases accessing this true amount of uncertainty is only ever possible with simulations. Below, we formularize the verification process for uncertainty calibration for both Bayesian and frequentist perspectives.
@@ 3.1.1 Bayesian Calibration
The accuracy of an approximator $q(\theta\mid y)$ of the true posterior $\pi(\theta\mid y)$ is difficult to study analytically, since the latter is rarely known in the first place. Fortunately, certain Bayesian self-consistency properties can be used even without knowledge about $\pi(\theta\mid y)$. Consider, for instance, a target quantity of interest $T=T(\theta)$, which may be derived from the parameters $\theta$. This could include the parameters themselves or any pushforward quantity, such as posterior predictions. For all target quantities $T$ and all uncertainty regions $U_{\alpha}(T(\theta)\mid y)$ obtained from $\pi(\theta\mid y)$ with nominal coverage probability $\alpha$ (e.g., credible intervals based on posterior quantiles), we know that, for the true value $T^{*}=T(\theta^{*})$, we have:
$\alpha=\int\int\mathbb{I}(T^{*}\in U_{\alpha}(T\mid y))\,\pi(y\mid\theta^{*})\,\pi(\theta^{*})\,\mathrm{d}y\,\mathrm{d}\theta^{*}.$ (1)
In other words, averaged over the data-generating process, the probability that an uncertainty region with coverage probability $\alpha$ contains the true value must be equal to $\alpha$. For a univariate quantity $T$, we can formulate this requirement as follows: under perfect Bayesian calibration, the posterior probability $p:=\Pr_{\pi(\theta\mid y)}(T\leq T^{*})$ (i.e., the “Bayesian $p$-value”) is uniformly distributed between 0 and 1 (Talts et al.,, 2018; Modrák et al.,, 2023; Gneiting et al.,, 2007). For a good approximation $q(\theta\mid y)$ of the true posterior $\pi(\theta\mid y)$, the above property should hold within tolerable error bounds. Unfortunately, the associated integrals in Equation 1 are not analytic, so none of their properties can be verified directly for $q(\theta\mid y)$. However, we can employ simulations from the underlying P model. For a data set $y$, we define
@@
where $\theta^{(m)}\sim q(\theta\mid y)$ are $M$ draws from the approximate posterior and $T^{(m)}=T(\theta^{(m)})$. If the posterior approximator is equal to the true posterior, the distribution of the posterior probabilities $p(T^{*},T^{(1:M)}\mid y)$ will follow a discrete uniform distribution between 0 and 1 (Modrák et al.,, 2023).
To test this property empirically, we first sample $S$ parameter draws from the prior, ${\theta^{*}}^{(s)}\sim\pi(\theta^{*})$ and subsequently $S$ data sets $y^{(s)}\sim\pi(y\mid{\theta^{*}}^{(s)})$ from the likelihood. Then, we obtain $M$ draws $\theta^{(m)}\sim q(\theta\mid y^{(s)})$ from the approximate posterior to compute $p^{(s)}:=p(T^{*},T^{(1:M)}\mid y^{(s)})$. Finally, we can test the set of posterior probabilities $\{p^{(s)}\}$ for uniformity (see Figure 3 for an illustration). Together, this simulation-based procedure is known as simulation-based calibration (SBC) checking (Talts et al.,, 2018; Modrák et al.,, 2023). SBC requires nested simulation: in the outer loop, we simulate parameters ${\theta^{*}}^{(s)}$ and data $y^{(s)}$ from the joint model $\pi(\theta^{*})\,\pi(y\mid\theta^{*})$ and in the inner loop, we sample $\theta^{(m)}$ from the approximate posterior $q(\theta\mid y^{(s)})$. As a result, SBC is typically computationally intensive, yet it remains a powerful and scale-free method for validating approximate posteriors.
@@ 3.1.2 Frequentist Calibration
Frequentist calibration works very similarly to Bayesian calibration, with two main differences. First, we do not sample true parameter values from a prior, $\theta^{*}\sim\pi(\theta^{*})$, but rather treat $\theta^{*}$ as fixed at some value. Second, our target of inference is not the true posterior $\pi(\theta\mid y)$, but a point estimator $\hat{\theta}=\hat{\theta}(y)$ of the parameters, along with the true sampling distribution $\pi(\hat{\theta}\mid\theta^{*})$ of the point estimator. This, in turn, implies a point estimator $\hat{T}=T(\hat{\theta})$ for any quantity of interest $T$, with a corresponding true sampling distribution $\pi(\hat{T}\mid T^{*})$. The latter can be transformed into uncertainty regions $U_{\alpha}(\hat{T}\mid y)$ (i.e., confidence intervals), which, under perfect frequentist calibration, will contain the true value $T^{*}$ for $\alpha$ percent of the datasets generated from $\pi(y\mid\theta^{*})$:
@@ 3.1.3 Power Analysis
The frequentist self-consistency of $p$-values holds only if the true parameter value $\theta^{*}$ used for simulating data from the likelihood $\pi(y\mid\theta^{*})$ is the same parameter assumed in the sampling distribution $\pi(\hat{\theta}\mid\theta^{*})$. In a null-hypothesis significance framework (see also Section 4.1), we consider a sampling distribution $\pi(\hat{\theta}\mid\theta_{0})$ given the parameter value $\theta_{0}$ characterizing the null-hypothesis. If the $\theta^{*}$ assumed in the sampling process $\pi(y\mid\theta^{*})$ is different from $\theta_{0}$, uniformity of the $p$-value distribution is no longer expected. Instead, when $\theta^{*}$ represents an explicit alternative hypothesis, the $p$-value distribution measures statistical power. Simulations are the prime tool for performing power analysis since analytical expressions of sampling distribution are often only available for few special parameter values, usually representing the null-hypothesis (Cohen,, 2013). Power analysis, evaluated via simulations, can also be performed in a Bayesian framework by choosing a different prior, $\pi(\theta_{0})$, within the estimated model than the prior, $\pi(\theta^{*})$, used in the data-generating process.
@@ 3.2 Other Simulation Studies
Good uncertainty calibration is undoubtedly an important property that should ideally be verified before entrusting the pair of probabilistic model and approximator with performing inference on real data. However, calibration alone is often insufficient. For example, when only testing parameters for calibration in a Bayesian model, $T(\theta)=\theta$, then not only the posterior $\pi(\theta\mid y)$ but also the prior $\pi(\theta)$ would pass the above described uniformity tests (Modrák et al.,, 2023). Accordingly, we have to additional study the sharpness of uncertainty estimators (Gneiting et al.,, 2007), which essentially indicates how small the uncertainty regions are for any fixed $\alpha$ (see also Bürkner et al.,, 2023). Among all well-calibrated uncertainty estimators, we would then choose the one that is sharpest. As for calibration, sharpness is almost always non-analytic so simulations are the only way to study it.
@@ 4 Model Inference
Simulations can be used also directly for model inference (Cranmer et al.,, 2020), that is, to estimate parameters or obtain decisions from the triple of statistical model, approximator, and observed data.
@@ 4.1 Hypothesis testing
Figure 4: Illustration of a simulation-based test. In the chosen scenario, the means of two groups are compared using a two-sample $t$-test for equal variances. Two independent data sets, each consisting of $N=40$ observations, were simulated from $\text{LogNormal}(\mu=2,\sigma=2)$. The null hypothesis equal group means holds and so does the equal variance assumption. However, the data deviate substantially from normality. Accordingly, even for $N=40$ per group, the left plot shows that the true distribution of the test statistics (approximated via simulations; red histogram), is clearly different from the large sample approximation assumed by the t-test (blue density). The right plot shows that the corresponding $p$-value distribution under the null hypothesis is highly non-uniform for the large sample approximation (blue histogram). In contrast, $p$-values obtained via a simulation-based test ($S=10000$) are almost perfectly uniform, with only small random error (red histogram).
@@
In some cases, the sampling distribution under the null hypothesis is indeed analytic, for instance, as in $t$-tests for normally distributed data, but often enough it is not. In such cases, we may use large-sample approximations, which indicate that for sufficiently large data sets, the test statistic’s distribution closely follows a known, tractable distribution (Ibragimov and Has’minskii,, 2013). However, the accuracy of these approximations for smaller sample sizes can be questionable (see Figure 4).
Simulation-based tests provide a powerful alternative to analytic solutions or large-sample approximations (Racine and Mackinnon,, 2007; Vasishth and Broe,, 2010). They generally proceed as follows: Sample $S$ draws $y_{0}^{(s)}\sim\pi(y\mid\theta_{0})$ from the likelihood given $\theta_{0}$. Then, compute the test statistic $T_{0}^{(s)}:=T(y_{0}^{(s)})$ for each draw. This implies $T_{0}^{(s)}\sim\pi(T(y)\mid\theta_{0})$, so we gained simple access to samples from the target distribution $\pi(T(y)\mid\theta_{0})$ (see Figure 4 left). Then, we compare the test statistic value $\tilde{T}:=T(y_{\text{obs}})$ from the observed data $y_{\text{obs}}$ with the distribution of simulated test statistic values $\{T_{0}^{(s)}\}$.
Suppose we are interested in a one-sided test for lower values of $T$. We would first compute $\tilde{r}=\sum_{s=1}^{S}\mathbb{I}(T_{0}^{(s)}<\tilde{T})$, where $\mathbb{I}$ is the indicator function. Then, we would compute the $p$-value $\tilde{p}$ corresponding to $y_{\text{obs}}$ as the normalized rank $\tilde{p}:=\tilde{r}/S$ (see Figure 4, right) and the significance threshold $T_{\rm\alpha}$ as the empirical $\alpha$-percentile of a set $\{T_{0}^{(s)}\}$ of simulated test statistics.
@@
Figure 5: Graphical illustration of frequentist vs. Bayesian approaches to simulation-based inference (SBI). Panel A: In frequentist SBI, a classifier is trained on simulated data and parameters $(y,\theta)$ to efficiently estimate a test statistic $\lambda(y;\theta)$ (e.g., the odds function). This statistic is used to compute critical and p-values via further simulations. An approximate confidence set for $\theta$ is constructed by inverting a series of hypothesis tests. The coverage of this set can be diagnosed through additional simulations (adapter after Dalmasso et al.,, 2024). Panel B: In Bayesian SBI, a generative neural network is trained to estimate the posterior $p(\theta\mid y)$ from simulated $(y,\theta)$ pairs. The calibration and sharpness of the resulting estimator can be evaluated via further simulations. Given observed data, the estimator can rapidly sample from the approximate posterior. Further simulations can be used to evaluate the fidelity of these samples via predictive checks or out-of-distribution (OOD) detection.

### v2

@@ 1 Introduction
In statistics, the emergence of simulation is typically associated with the birth of Monte Carlo methods during World War II (Robert and Casella,, 2011), eventually leading to the publication of the Metropolis algorithm (Metropolis et al.,, 1953). Hastings, (1970) and Peskun, (1973) later generalized the Metropolis algorithm as a family of simulation-based tools, namely, Markov chain Monte Carlo (MCMC), designed to mitigate the curse of dimensionality in direct Monte Carlo estimation. Crucially, the inception of Monte Carlo methods coincides with the end of what Efron and Hastie, (2021) characterize as classical statistical inference and the beginning of early computer-age methods, such as bootstrap and jackknife, that demand randomized resampling and mechanical repetition.
@@
Early MCMC methods freed Bayesian analysis from the confines of conjugate models and closed-form posteriors, most notably through the introduction of the Gibbs sampler (Geman and Geman,, 1984; Gelfand and Smith,, 1990) and hybrid variants (Tierney,, 1994). However, these methods still required explicit models with tractable likelihoods. In contrast, implicit models—defined through a stochastic mechanism (e.g., a randomized simulation) without a known distribution for computing the likelihood of their outputs (Diggle and Gratton,, 1984)—remained out of reach for Bayesian methods until the advent of approximate Bayesian computation (ABC; Rubin,, 1984; Tavaré et al.,, 1997). In ABC, stochastic simulation explicitly takes on a dual role: it serves both as an epistemic tool for solving forward problems (from unknowns to knowns) and as a computational tool for solving inverse problems (from knowns to unknowns). Indeed, the coinage of the term simulation-based inference (SBI; Cranmer et al.,, 2020) is reflective of this role.
Figure 1: Forward simulation and inverse inference are two central components in a statistical model of unknowns $\theta$ and observables $y$.
As the notion of statistical model loses its well-defined contours in an AI-dominated conceptual landscape, the role of simulation in statistics becomes increasingly difficult to delineate from its role in science in general. However, since this paper focuses on statistical workflows—iterative processes of model building and model criticism—our discussion will center on the versatility of simulating from statistical models. In particular, this means generating random draws from model-implied distributions, much like drawing random cards from a cleverly arranged deck. Given this focus, we will not detail sampling algorithms that primarily rely on evaluating the density of statistical models rather than their generative capabilities.
Among density-based approaches, the most prominent family of algorithms is undoubtedly MCMC (Brooks et al.,, 2011). MCMC has seen rapid developments in recent years, including locally adaptive methods that can handle distributions with complex geometries (Biron-Lattes et al.,, 2024; Modi et al.,, 2024) and methods that leverage modern hardware for large-scale parallelization (Margossian et al.,, 2024; Sountsov et al.,, 2024). Other popular families of algorithms are Sequential Monte Carlo (SMC; Del Moral et al.,, 2006) and importance sampling (Tokdar and Kass,, 2010). Even though these algorithms involve simulation steps, they typically do not require direct draws from the model’s generative distribution, but instead rely on simulating proposals or particles guided by density evaluations.
The organization of the following sections elucidates use cases for statistical model simulations throughout the key stages of a minimalist statistical workflow, namely: 1) model specification; 2) model verification; 3) model inference; and 4) model checking. We conclude with a brief outlook on simulation intelligence and the ascent of automated statistics.
@@ 2 Model Specification
  • P (Probability Distribution) – the probabilistic joint model (hereafter called P model) that defines the relationships between model parameters, latent variables, and observed data.
@@
According to the PAD model taxonomy, model specification refers to defining the components of a P model. In a Bayesian setting, this entails formulating a prior $\pi(\theta)$ and an observation model $\pi(y\mid\theta)$ which together determine the generative joint model $\pi(\theta,y)$. In a frequentist setting, the P model reduces to the observation model $\pi(y\mid\theta)$, which is defined over a specific domain $\Theta\ni\theta$.
As part of a Bayesian workflow (Gelman et al.,, 2020), the plausibility of P models can already be visualized and evaluated through prior predictive or prior pushforward checks (Gabry et al.,, 2019). The purpose of these checks is to assess whether the generative behavior of a P model is consistent with the available domain expertise via simulations from the model-implied (i.e., prior predictive) distribution, $\pi(y)=\int\pi(y\mid\theta)\pi(\theta)\mathrm{d}\theta$. Consistency with domain expertise (i.e., known knowns and known unknowns) not only figures in domain-specific Bayesian workflows (Schad et al.,, 2021), but can also be considered as a crucial aspect of iterative theory building in general.
Here, once again, simulation serves as an antidote to complexity: consistency with domain expertise may not be immediately evident from the model formulation alone, even when a model’s specification appears beguilingly simple (see, for example, Aguilera-Venegas et al.,, 2019, for a probabilistic extension of Conway’s Game of Life). Moreover, if the data $y$ is high-dimensional, the utility of consistency checks can be challenging to assess. This is where prior pushforward checks come into play (Gabry et al.,, 2019), following the procedure suggested by Schad et al., (2021): (1) define a low-dimensional, interpretable summary statistic $T(y)$; (2) determine plausible value regions for the summary statistic; and (3) simulate model predictions to ensure the summary statistic falls within these regions [^1] In a frequentist setting, the prior can be replaced with a set of plausible values for the parameters, $\Theta_{\text{plausible}}$..
@@
Figure 2: Graphical illustration of the simulation-based framework for prior elicitation proposed by Bockting et al., 2024b . The process starts by identifying target quantities via the domain expert and eliciting expert statistics $\hat{T}$. Model predictions are then simulated by sampling from a parametric prior $\pi(\theta\mid\lambda)$ and computing model-implied target quantities $T$. A loss function $L$ assesses the consistency between model and expert-elicited statistics and adjusts prior hyperparameters $\lambda$ to minimize discrepancies. The process continues until prior predictions align closely with expert knowledge.
@@ 3 Model Verification
Before deploying a statistical model on real data, we should first ensure that the interplay between the probabilistic joint model P and the approximator A functions as intended in the scenarios they were designed for, namely, when their underlying assumptions hold. In many cases, simulations offer the only feasible means of verifying such correctness, or at least gauging in silico whether inferences are sufficiently informative for their intended purposes.
@@ 3.1 Verifying Calibration
Proper uncertainty calibration is is crucial for ensuring confidence in estimates and predictions from statistical models (Little,, 2006; Gneiting et al.,, 2007). Informally, we can say that the uncertainty of an approximator is well calibrated if it captures the true amount of uncertainty in a system. Except for special cases, accessing this true uncertainty is only ever possible with simulations. Below, we formalize the verification process for uncertainty calibration for both Bayesian and frequentist perspectives.
@@ 3.1.1 Bayesian Calibration
In the following, we will denote an approximator of the true posterior $\pi(\theta\mid y)$ as $q(\theta\mid y)$. This $q(\theta\mid y)$ may be given implicitly, for example, in the form of an MCMC algorithm or explicitly, for example, in the form of a generative neural network. The accuracy of an approximator $q(\theta\mid y)$ is difficult to study analytically, since the true posterior $\pi(\theta\mid y)$ is rarely known in the first place. Fortunately, certain Bayesian self-consistency properties can be used even without knowledge about $\pi(\theta\mid y)$. Consider, for instance, a target quantity of interest $T=T(\theta)$, which may be derived from the parameters $\theta$. This could include the parameters themselves or any pushforward quantity, such as posterior predictions. For all target quantities $T$ and all uncertainty regions $U_{\alpha}(T(\theta)\mid y)$ obtained from $\pi(\theta\mid y)$ with nominal coverage probability $\alpha$ (e.g., credible intervals based on posterior quantiles), we know that
$\alpha=\int\int\mathbb{I}(T^{*}\in U_{\alpha}(T\mid y))\,\pi(y\mid\theta^{*})\,\pi(\theta^{*})\,\mathrm{d}y\,\mathrm{d}\theta^{*},$ (1)
where $\theta^{*}\sim\pi(\theta^{*})$ are prior draws that serve as the ground truth for the corresponding target quantity $T^{*}=T(\theta^{*})$. In other words, the probability that an uncertainty region with coverage probability $\alpha$ contains the true value must be equal to $\alpha$ on average over the data-generating process. For a univariate quantity $T$, we can formulate this requirement as follows: under perfect Bayesian calibration, the posterior probability $p:=\Pr_{\pi(\theta\mid y)}(T\leq T^{*})$ (i.e., the “Bayesian $p$-value”) is uniformly distributed between 0 and 1 (Cook et al.,, 2006; Gneiting et al.,, 2007; Talts et al.,, 2018). For a good approximation $q(\theta\mid y)$ of the true posterior $\pi(\theta\mid y)$, the above property should hold within tolerable error bounds. Unfortunately, the associated integrals in Equation 1 are not analytic, so none of their properties can be verified directly for $q(\theta\mid y)$. However, we can employ simulations from the underlying P model. For a data set $y$, we define
@@
where $\theta^{(m)}\sim q(\theta\mid y)$ are $M$ draws from the approximate posterior and $T^{(m)}=T(\theta^{(m)})$. If the posterior approximator is equal to the true posterior, the distribution of the posterior probabilities $p(T^{*},T^{(1:M)}\mid y)$ will follow a discrete uniform distribution between 0 and 1 (Modrák et al.,, 2025).
To test this property empirically, we first sample $S$ parameter draws from the prior, ${\theta^{*}}^{(s)}\sim\pi(\theta^{*})$ and subsequently $S$ data sets $y^{(s)}\sim\pi(y\mid{\theta^{*}}^{(s)})$ from the likelihood. Then, we obtain $M$ draws $\theta^{(m)}\sim q(\theta\mid y^{(s)})$ from the approximate posterior to compute $p^{(s)}:=p(T^{*},T^{(1:M)}\mid y^{(s)})$. Finally, we can test the set of posterior probabilities $\{p^{(s)}\}$ for uniformity (see Figure 3 for an illustration). More powerful methods for diagnosing joint calibration are possible, including training a classifier to learn test statistics directly from simulations (Yao and Domke,, 2023; Bansal et al.,, 2025) or using the model likelihood if available (Modrák et al.,, 2025).
These procedures for checking Bayesian calibration are collectively referred to as simulation-based calibration (SBC) (Talts et al.,, 2018; Modrák et al.,, 2025). Crucially, SBC requires nested simulations: in the outer loop, we simulate parameters ${\theta^{*}}^{(s)}$ and data $y^{(s)}$ from the joint model $\pi(\theta^{*})\,\pi(y\mid\theta^{*})$; in the inner loop, we sample $\theta^{(m)}$ from the approximate posterior $q(\theta\mid y^{(s)})$. As a result, SBC is typically computationally intensive, yet it remains a general method for validating approximate posteriors.
@@ 3.1.2 Frequentist Calibration
[…]
@@ 3.1.3 Power Analysis
The frequentist self-consistency of $p$-values holds only if the true parameter value $\theta^{*}$ used for simulating data from the likelihood $\pi(y\mid\theta^{*})$ is the same parameter assumed in the sampling distribution $\pi(\hat{\theta}\mid\theta^{*})$. In a null-hypothesis significance framework (see also Section 4.1), we consider a sampling distribution $\pi(\hat{\theta}\mid\theta_{0})$ given the parameter value $\theta_{0}$ characterizing the null-hypothesis. If the $\theta^{*}$ assumed in the sampling process $\pi(y\mid\theta^{*})$ is different from $\theta_{0}$, uniformity of the $p$-value distribution is no longer expected. Instead, when $\theta^{*}$ represents an explicit alternative hypothesis, the $p$-value distribution measures statistical power. Simulations are the prime tool for performing power analysis since analytical expressions of sampling distribution are often only available for few special parameter values, usually representing the null-hypothesis (Cohen,, 2013). Simulation-based power analysis can also be performed in a Bayesian framework by choosing a prior $\pi(\theta_{0})$ within the estimated model that differs from the prior $\pi(\theta^{*})$ used in the data-generating process.
@@ 3.2 Other Simulation Studies
Good uncertainty calibration is undoubtedly an important property that should ideally be verified before entrusting the pair of probabilistic model and approximator with performing inference on real data. However, calibration alone is often insufficient. For example, when only testing parameters for calibration in a Bayesian model, $T(\theta)=\theta$, then not only the posterior $\pi(\theta\mid y)$ but also the prior $\pi(\theta)$ would pass the above described uniformity tests (Modrák et al.,, 2025). Accordingly, we have to additionally study the sharpness of uncertainty estimators (Gneiting et al.,, 2007), which essentially indicates how small the uncertainty regions are for any fixed $\alpha$ (see also Bürkner et al.,, 2023). Among all well-calibrated uncertainty estimators, we would then choose the one that is sharpest. As for calibration, sharpness is almost always non-analytic so simulations are the only way to study it.
@@ 4 Model Inference
Simulations can also be used directly for model inference (Cranmer et al.,, 2020), that is, to estimate parameters or obtain decisions from the triple of statistical model, approximator, and observed data.
@@ 4.1 Hypothesis testing
Figure 4: Illustration of a simulation-based test. In the chosen scenario, the means of two groups are compared using a two-sample $t$-test for equal variances. Two independent data sets, each consisting of $N=40$ observations, were simulated from $\text{LogNormal}(\mu=2,\sigma=2)$. The null hypothesis asserts equal group means and equal variances. However, the data deviate substantially from normality. Accordingly, even for $N=40$ per group, the left plot shows that the true distribution of the test statistics (approximated via simulations; red histogram), is clearly different from the large sample approximation assumed by the t-test (blue density). The right plot shows that the corresponding $p$-value distribution under the null hypothesis is highly non-uniform for the large sample approximation (blue histogram). In contrast, $p$-values obtained via a simulation-based test ($S=10,000$) are almost perfectly uniform, with only small random error (red histogram).
@@
In some cases, the sampling distribution under the null hypothesis is indeed analytic, for instance, as in $t$-tests for normally distributed data, but often enough it is not. We may then appeal to asymptotic approximations, under which the (sampling) distribution of the test statistic is a well-characterized distribution in the limit of an infinite number of simulations (Ibragimov and Has’minskii,, 2013). However, the accuracy of these approximations for smaller sample sizes can be questionable (see Figure 4).
Simulation-based tests provide a powerful alternative to analytic solutions or large-sample approximations (Racine and Mackinnon,, 2007; Vasishth and Broe,, 2010). They generally proceed as follows: Sample $S$ draws $y_{0}^{(s)}\sim\pi(y\mid\theta_{0})$ from the likelihood given $\theta_{0}$. Then, compute the test statistic $T_{0}^{(s)}:=T(y_{0}^{(s)})$ for each draw, which implies $T_{0}^{(s)}\sim\pi(T(y)\mid\theta_{0})$. Thereby, we gain access to samples from the target distribution $\pi(T(y)\mid\theta_{0})$ simply via predefined transformations of individual samples (see Figure 4 left). Then, we compare the test statistic value $\tilde{T}:=T(y_{\text{obs}})$ from the observed data $y_{\text{obs}}$ with the distribution of simulated test statistic values $\{T_{0}^{(s)}\}$.
[…]
@@
Figure 5: Graphical illustration of frequentist vs. Bayesian approaches to simulation-based inference (SBI). Panel A: In frequentist SBI, a classifier is trained on simulated data and parameters $(y,\theta)$ to efficiently estimate a test statistic $\lambda(y;\theta)$ (e.g., the odds function). This statistic is used to compute critical and p-values via further simulations. An approximate confidence set for $\theta$ is constructed by inverting a series of hypothesis tests. The coverage of this set can be diagnosed through additional simulations (adapter after Dalmasso et al.,, 2024). Panel B: In Bayesian SBI, a generative neural network is trained to estimate the posterior $\pi(\theta\mid y)$ from simulated $(y,\theta)$ pairs. The calibration and sharpness of the resulting estimator can be evaluated via further simulations. Given observed data, the estimator can rapidly sample from the approximate posterior. Further simulations can be used to evaluate the fidelity of these samples via predictive checks or out-of-distribution (OOD) detection.